OpenStudy (anonymous):

how would you rewrite 8% as a decimal number?? or any number so i know how to do it

OpenStudy (anonymous):

would it be .8?

OpenStudy (anonymous):

0.8?

OpenStudy (anonymous):

No. divide 8 by 100 to change from a percent to a decimal. 0.08

OpenStudy (anonymous):

so for any number you would divide by 100? unless its over 100 then you'd divide by 1,000? @DSS

OpenStudy (petiteme):

0.08

OpenStudy (anonymous):

no...to change every percent, you divide by 100

OpenStudy (anonymous):

125% would become 125/100 or 1.25

OpenStudy (anonymous):

100% = 1.0 in decimal form it means 1 whole i.e. 100% i.e all. 100% = 100/100 = 1

OpenStudy (anonymous):

okay thankyou!

OpenStudy (anonymous):

8 %=8/100=.08